OpenStudy (anonymous):

let \[f(x) = \sqrt[3]{x^2+4x}\] and let g(x) be antiderivative of f(x). then if g(5) = 7 then find g(1)

OpenStudy (anonymous):

let \[f(x) = \sqrt[3]{x^2+4x}\] and let g(x) be antiderivative of f(x). then if g(5) = 7 then find g(1)

OpenStudy (amistre64):

rutn it into an exponenetial form

OpenStudy (anonymous):

hmm

OpenStudy (amistre64):

\[f(x) = ({x^2+4x})^{1/3}\]

OpenStudy (amistre64):

turn it into

OpenStudy (amistre64):

then power rule it, but always include the chain rule that pops out with it

OpenStudy (anonymous):

hmm k give me a sec

OpenStudy (amistre64):

\[f(x) = ({x^2+4x})^{1/3}\] \[g(x) = \frac{1}{3}({x^2+4x})^{-2/3}\ *[x^2+4x]'\] \[ g(x)= \frac{1}{3}({x^2+4x})^{-2/3}\ (2x+4)\] \[ g(x)= \frac{2x+4}{3(x^2+4x)^{2/3}}\] but thats just me deriving it .... you want the integration ...

OpenStudy (anonymous):

what about it : g '(x) = f(x) and then \[g(x) = \int\limits_{a}^{x}\sqrt[3]{t^2 + 4t}dt\] then we let h(x) = g(x) - 7 and \[h(x) = \int\limits_{5}^{x}\sqrt[3]{t^2 + 4t}dt\]

OpenStudy (anonymous):

yeap i want the integration :D

OpenStudy (amistre64):

maybe integration by parts?

OpenStudy (anonymous):

mybe :) so you gonna do that? \[h(r) = \int\limits_{5}^{x}\sqrt[3]{t^2}dt + \int\limits_{5}^{x}\sqrt[3]{4t}dt\]

OpenStudy (amistre64):

lol .... maybe not http://www.wolframalpha.com/input/?i=int+%28x^2%2B4x%29^%281%2F3%29+dx

OpenStudy (amistre64):

turn it into a power series if need be .... and play with the poly that you get

OpenStudy (amistre64):

if you think its gonna be an esay split; it aint, not if the wolf spits back that thing

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

i think i just did a simple mistake on using integration by parts o.O , i forgot a rule i guess you can't use it in roots! ??

OpenStudy (anonymous):

this integral sucks. there must be some other gimmick at work here to solve the problem without using integration

OpenStudy (anonymous):

o.O hmm

OpenStudy (anonymous):

i see no snap way to answer this question. maybe i am missing something

OpenStudy (anonymous):

well i think my way solves the question mybe? last i said that \[h(x) = \int\limits_{5}^{x}\sqrt[3]{t^2 + 4t}\] and if we will solve it for h(1) (h(x) is an antiderivative of f(x) .... we solve that and we got (calculator <3 <3 ) -10.88222 since h(x) = g(x) - 7 , g(x) = h(x) + 7 and i have to go to dinner brb

OpenStudy (anonymous):

at the moment( h(x) = g(x) - 7 )

OpenStudy (anonymous):

oh ok now back give me a minute

OpenStudy (anonymous):

and it fallows that : g(x) = h(x) + 7 then we can do that g(1) = h(1) + 7 = -10.88222 + 7 = -3.88222 !!!!!!!!!!

OpenStudy (anonymous):

err is this true what i did?

OpenStudy (anonymous):

so if we look what we did : take g(x) as antiderivative of f(x) than we will get g'(x) = f(x) or \[g(x) = \int\limits_{a}^{x}\sqrt[3]{t^2 + 4t}dt\] but its unsolveable so instead of integrating we let p(x) = g(x) - 7 (p(x) is also an antiderivative of f(x) and thann we can write \[p(x) = \int\limits_{5}^{x} \sqrt[3]{t^2 + 4t}\] !!! now by using calculator we easly got \[p(1) = \int\limits\limits_{5}^{1} \sqrt[3]{t^2 + 4t}\] = 10.88222 (CALCULATOOOR <3!!) we said p(x) = g(x) - 7 g(x) = p(x) + 7 so that g(1) = p(1) + 7 = -10.88222 + 7 = -3.88222

OpenStudy (anonymous):

i guess its right :D

OpenStudy (zarkon):

\[\int\limits_{5}^{x}f(t)dt=g(x)-g(5)\] \[g(x)=\int\limits_{5}^{x}f(t)dt+7\] so \[g(1)=\int\limits_{5}^{1}f(t)dt+7=-\int\limits_{1}^{5}f(t)dt+7\approx -3.88222\]

OpenStudy (anonymous):

hold the phone where did the number come from???

OpenStudy (anonymous):

which number? you mean limits?

OpenStudy (anonymous):

no the 10.333 whatever?

OpenStudy (anonymous):

-10.88 etc

OpenStudy (anonymous):

use calculation to solve the integral and you get that result :D

OpenStudy (anonymous):

ooooooooooooooh. i thought it was some theory you were supposed to use! numerical integration. ok

OpenStudy (anonymous):

and i used second fundemental theorem of calculus here :D

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

lol :)

OpenStudy (anonymous):

you used ti 89, wolfram, whatever

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

yeap i just made the integral as solvable, found limits :D used second fundemental theorem of calculus than used a calculator to solve the integral :D :D

OpenStudy (anonymous):

hello! integrate from 1 to 5 and adjust. cheating if you ask me...

OpenStudy (anonymous):

Ahaha man com'on :D its the solution anyway

OpenStudy (anonymous):

hi imranmeah!

OpenStudy (anonymous):

no i like it it is good for sure

OpenStudy (anonymous):

3 mounths ago, i did not even know what limits are :P

OpenStudy (anonymous):

(i mean i did not kno derivatives, lineer algebra, 0 knowledge of algebra and calculus) :D

OpenStudy (zarkon):

you can get bounds for the solution by using the mean value theorem...But I see no nice way of getting the exact answer.

OpenStudy (anonymous):

u right, but i'm not sure of it

OpenStudy (anonymous):

that is for sure!